OpenStudy (gabbyalicorn):

A circular swimming pool has a diameter of 12 m. What is the approximate circumference of the swimming pool? Use π= 3.14. Express your final answer to the nearest hundredth.

OpenStudy (gabbyalicorn):

I'm thinking 37.68. \[3.14 \]

OpenStudy (gabbyalicorn):

3.14 x 12

OpenStudy (gabbyalicorn):

@UsukiDoll @unknownunknown

OpenStudy (usukidoll):

C = 2pir? \[C = 2 \pi r \]

OpenStudy (usukidoll):

my head hurts too much thinking

OpenStudy (gabbyalicorn):

oh sorry

OpenStudy (gabbyalicorn):

what is the radius?

OpenStudy (gabbyalicorn):

@johnweldon1993

OpenStudy (johnweldon1993):

Remember \(\large Radius = \frac{1}{2}Diameter\)

OpenStudy (gabbyalicorn):

so 2 x 3.14 x 2?

OpenStudy (johnweldon1993):

But remember the equation? \[\large C = 2\pi r\] we can also write that as \[\large C = \pi d\] right? *2r = d* So you were right on the money before...12 times 3.14 = ?

OpenStudy (gabbyalicorn):

37.68

OpenStudy (gabbyalicorn):

:D okay, thanks!

OpenStudy (johnweldon1993):

Right :)